allcm(1) teTeX allcm(1) NAME allneeded−forcethecalculationofallfontsnowneeded SYNOPSIS allneeded[−r]files... DESCRIPTION allneededforcesthecalculationofallfontsthatareneededtopreviewasetofdvifiles.Justspecifywheretheprogramshouldsearchforfilesonthecommandline.Thefontsgenerationistriggeredbyrunningdvips(1)overallaccessibledvi−files.Therefore,thefontsarecreatedintheresolutionneededbydvips. Ifthe−rflagisspecified,thecommanddvired(1)willbeusedinsteadofdvips. allneededdoesnotrecalculateexistingfonts(aslongastheMetafontmodedoesnotchange). SEEALSO alldc(1),allcm(1),dvips(1),dvired(1),fontimport(1),fontexport(1). AUTHOR DirkHillbrecht(dh@informatik.uni−hannover.de) teTeX 11/94 1